# with application to sea and land ice

### v1

@@ Introduction
The framework presented in section 2 suggests an alternative approach to rheology inference: Instead of learning the effective shear viscosity $\psi$, learn the potential of dissipation $j$. The potential $j$ is introduced in section 2.1 to establish conditions under which the governing equations are well-posed. To this end, we enforce the convexity of $j$ or, equivalently, the monotonicity of $s\mapsto\psi(s,\lambda)s$ when the fluid is one-dimensional or incompressible. Parameterizing the convex potential $j$ with input convex NNs, which have demonstrated accurate interpolation performance for convex models [amos2017], would remove the need for the monotonicity penalty. Moreover, for two-dimensional compressible fluids, this approach ensures the existence of a convex potential of dissipation (in two dimensions, unlike the one-dimensional or incompressible fluid case, an underlying potential of dissipation does not exist in general). Input convex NNs are used in [parolini2025] to represent convex effective viscosities successfully. Despite many rheological models being convex, this is not always the case; see e.g. the effective shear viscosities of polymer solutions in [laun1984].

### v2

@@ Introduction
The framework presented in section 2 suggests an alternative approach to rheology inference: Instead of learning the effective shear viscosity $\psi$, learn the potential of dissipation $j$. The potential $j$ is introduced in section 2.1 to establish conditions under which the governing equations are well-posed. To this end, we enforce the convexity of $j$ or, equivalently, the monotonicity of $s\mapsto\psi(s,\lambda)s$ when the fluid is one-dimensional or incompressible. Parameterizing the convex potential $j$ with input convex NNs, which have demonstrated accurate interpolation performance for convex models [3], would remove the need for the monotonicity penalty. Moreover, for two-dimensional compressible fluids, this approach ensures the existence of a convex potential of dissipation (in two dimensions, unlike the one-dimensional or incompressible fluid case, an underlying potential of dissipation does not exist in general). Input convex NNs are used in [34] to represent convex effective viscosities successfully. Despite many rheological models being convex, this is not always the case; see e.g. the effective shear viscosities of polymer solutions in [27].
@@ Acknowledgement
Data availability
The data and scripts required to reproduce the results in this article are contained in the following repository: https://bitbucket.org/gonzalogddiego/rheologyinference.